### v1

# Universal neighborhood topology and geometry of exceptional points in physical systems

Abstract
Exceptional points (EPs) occurring in non-Hermitian systems at certain physical parameters are intensively studied in many areas of physics, including diffraction optics, lasers, atomic and polaritonic condensates, often in the context of sensing. Recent discoveries of EPs in nonlinear systems open the door for an even larger parameter space, raising the question of whether the geometric structure of EPs is universal and independent of the physical model. We show that this is the case for nonlinear perturbations of an isolated 2nd-order linear EP which becomes the organizing point of a universal elementary catastrophe (elliptic umbilic). This clarifies not only the neighborhood’s topology but also its geometric shape (cone with quasi-deltoid cross section). Thus, the position and characteristics of EPs can be predicted in nonlinear non-Hermitian parameter space; e.g., at a 2nd-order linear EP four nonlinear eigenvectors coalesce. These fundamental insights on universal topological structures and phase boundaries accompanying EPs in nonlinear physical systems will pave the way for the purposeful design of such systems with novel functionalities and control possibilities.
Over the last decade or so, the physics of exceptional points (EPs), which are commonly defined as non-Hermitian degeneracies with coalescing eigenvalues and eigenvectors (mathematical definitions and generalizations are in kato.1966 ; ashida-etal.2021 ), has been increasingly explored across a broad range of scientific areas, including anharmonic oscillators bender-wu.1969 , diffraction optics berry-odell.1998 , quantum chaos leyvraz-heiss.2005 ; heiss.2012 , conventional lasers brandstetter-etal.2014 ; lietzer-etal.2012 , waveguides and photonic systems kullig-etal.2023 ; longhi.2018 ; quirozjuarez.2019 ; dong-etal.2020 , atomic gases, condensates and high-Q cavity atomic lasers xu-etal.2017 ; choi-etal.10 ; wang-etal.2024 , polariton lasers, microlasers and polariton condensates dembowski-etal.01 ; miao-etal.2016 ; gao-etal.15 ; gao2018chiral ; ozturk-etal.2021 ; khurgin.2020 ; Li2022 ; PhysRevLett.122.185301 ; PhysRevB.104.235408 ; Li2022 , non-Hermitian Bose–Hubbard models graefe-etal.2008 , critical fluctuations and fluctuation spectra hanai-littlewood.20 ; binder-kwong.2021 , and sensing applications wiersig.2014 ; hodaei-etal.2017 ; langbein.2018 ; wiersig.2020 ; sakhdari-etal.2019 , non-Hermitian parity-time (PT) symmetric quantum systems bender-hook.2024 ; hassan-etal.2015 ; miri-alu.2016 ; ge-elganainy.2016 ; kominis-etal.2017 ; teimourpour-etal.2017 ; xia-etal.2021 ; lee-etal.2024 ; klauck-etal.2025 , and quantum state control abbasi-etal.2022 . EPs have also been studied in the context of topological insulators and other topological systems (e.g. luitz-etal.2019 ; kawabata-etal.2019 ; tang-etal.2020 ; bergholtz-etal.2021 ; hu-etal.2022 ; yokomizo-etal.2020 ; ding-etal.2022 ). While in most of these physical realizations EPs emerge as eigenvalues $\lambda$ of a linear non-Hermitian matrix $M$, i.e. from the eigenvalue problem $M\textbf{x}=\lambda\textbf{x}$, more recent studies of EPs include nonlinear physical systems, such as optical or electric resonators, as well as polariton condensates. In these systems, the complex interplay of nonlinear and non-Hermitian physics shifts PhysRevA.103.043510 ; el2023tracking ; Ramezanpour:24 ; wingenbach-etal.2024 and rotates wingenbach-etal.2024 the EP in parameter space by blueshifts and saturable gain. Moreover, mode switching can be realized in bistable regions near the EP wang2019dynamics . In Refs. wingenbach-etal.2024 ; Li2024 , it was shown that nonlinearity can alter the eigenvalue splitting in the vicinity of an EP and highly sensitive EP-based sensors have been realized in nonlinear non-Hermitian systems PhysRevApplied.18.054059 ; Bai-etal.2023 .
The eigenvalue problem may be nonlinear in the sense that the matrix $M$ may depend on the eigenvector x. For example, the energies in an optical microcavity, which are described by the diagonal elements of $M$, can shift with the intensity of the light field described by x, hence $M(\textbf{x})\textbf{x}=\lambda\textbf{x}$. In the following, we refer to EPs that are solutions of a linear eigenvalue problem as ‘linear EPs,’ and those solving a nonlinear eigenvalue problem are called ‘nonlinear EPs.’
The wide range of physical systems in which EPs can occur raises the question whether there exist any guiding principles that help clarify the occurrence of EPs as the physical parameters entering the matrix $M$ in each of the systems are varied, for example, the mode coupling in classical diffraction gratings berry-odell.1998 or frequency and coupling constants of ring resonators in sensing applications wiersig.2020 ).
In this paper, we demonstrate that general insight can be obtained on the distribution of EPs in parameter space, especially for nonlinear eigenvalue problems. We show that a powerful way to do this is to consider the behavior of eigenvectors around EPs as a bifurcation problem and make use of established mathematical results in catastrophe theory. We illustrate this procedure with an analysis of (nonlinear) parameter-dependent $2\times 2$ matrices and show that, for a broad class of physically relevant nonlinear eigenvalue problems, there are general constraints to the topology and geometric shape of the set of nonlinear EPs. These constraints are universal and independent of the physical system that the matrix $M$ represents.
An EP occurs at specific parameter values. Let us assume the linear EP to be isolated, occurring at certain values of the parameters entering the linear matrix $M$. We can vary the parameters away from the EP, in other words look at the neighborhood of the EP, and, by assumption, we will not find any other linear EPs in a small neighborhood. But if we allow the matrix to be nonlinear, the question about EPs in the neighborhood becomes more complicated. To answer the question, we would have to solve the nonlinear eigenvalue problem in order to find the location of the EP – or possibly even a multitude of EPs. We could then see if, for a given nonlinearity, we have more EPs or if the EP goes away once the nonlinearity is switched on. But in order to get a complete picture of the EPs existing in the nonlinear case, we would have to fix the nonlinearity and then solve the nonlinear eigenvalue problem for all remaining parameters entering the matrix. It is a priory not even clear how many remaining parameters we have to take into consideration. For example, in the case of the coupled ring resonators, we could limit ourselves to cases without loss and with each resonator having the same frequency (3 parameters, cf. Eq. (3) in wiersig.2020 ), but we could also generalize the model if we assume each resonator to have a different frequency and different loss (5 parameters). And we would have to do this ambitious analysis for each fixed nonlinearity. Finally, we would have to carry out this program for each physical realization of EPs (diffraction optics, polariton physics, sensing applications, etc.) separately. This means that, realistically speaking, it is difficult to get a full understanding of the structure of the set of EPs.
To solve the highly sophisticated problem outlined above, we utilize the general theorems of a branch of bifurcation theory called catastrophe theory zeeman.1976b ; thom.1989 ; arnold.1992 ; saunders.80 ; gilmore.81 ; poston-stewart.1978 . Roughly speaking, the framework of catastrophe theory describes the shapes of singularity sets – or phase boundaries in nonequilibrium phase diagrams, to use a physicist’s language – related to singularities of potential functions $V_{L}$ (Lyapunov potentials). The word catastrophe is used for the sudden qualitative change or bifurcation at a ‘phase boundary.’ Moreover, it lists all possible shapes (‘elementary catastrophes’) for cases in which the number of variables (often called state variables) and the number of control parameters (i.e. external parameters) are relatively small, see Supplemental Material (SM), Table 1).
Examples of systems that have been analyzed by catastrophe theory include dynamical systems obeying equations of motion of the form $=-{\rm grad}\{V_{L}(\textbf{x})\}$ x ) }, where the phase boundaries in the control parameter space follow from the assumption of degenerate stationary states (degenerate critical points), corresponding to ${\rm grad}\{V_{L}(\textbf{x})\}=0$ and ${\rm det}\{h\}=0$, where $h$ denotes the Hessian matrix of $V_{L}$. Similarly, in optics, the solution of the Helmholtz equation can be evaluated with the method of stationary phase and stationary points, as well as caustics occuring when two or more stationary points coalesce (p. 336 of adam.2002 ). This identifies rainbows and other optical phenomena as diffraction catastrophe, specifically fold and cusp bifurcations; see berry-upstill.1980 ; berry.90 ; adam.2002 and Ch. 12 of poston-stewart.1978 for details. Higher bifurcations (catastrophes), such as the so-called elliptic umbilic, have also been studied in the context of optical diffraction berry-etal.79 ; berry.90 . Many other areas of application of catastrophe theory can be found in zeeman.1976a ; zeeman.1976b ; saunders.80 ; gilmore.81 ; rosen.1988 ; poston-stewart.1978 ; heiss.04 ; longhi.2018 ; heiss.2012 ; barkleyrosser.2007 . This further extends to studies in biology, social sciences, and economics, with various degrees of acceptance because low-dimensional dynamical-systems models are not always appropriate for the simulation of complex systems in those sciences rosen.1988 ; barkleyrosser.2007 . In contrast, in many areas of physics low-dimensional models and eigenvalue problems (e.g. two–level models) are well established.
Figure 1: (a) Sketch of a non-Hermitian and nonlinear two-mode system, with mode coupling $\beta$, mode energy splitting $2\delta=\delta_{x}-\delta_{y}$, and loss/gain difference $2\gamma$. (b) Sketch of hypothetical location of exceptional points as a function of physical parameters. Without a nonlinearity, $\alpha=0$, isolated linear EPs are at $(\beta,\delta)=(\pm\gamma,0)$; in the nonlinear regime, $\alpha\neq 0$, EPs are implied on the colored surfaces. To illustrate the problem our study is solving, the surfaces we show here are one of infinitely many hypothetically possible (but generally wrong) EP neighborhoods. The correct geometric shape is shown in Fig. 2. Panels (c) and (d) show the eigenvalues in the linear case ($\alpha=0$) with two isolated EPs at $(\beta,\delta)=(\pm\gamma,0)$.
Figure 2: Surface of EP locations as a function of physical parameters (singularity set) explained in Fig. 1. In the weakly nonlinear regime, $\alpha\ll\gamma$, the shape approaches that of an elliptic umbilic bifurcation (‘catastrophe’). The shape of each singularity set consists of two infinite three–cusped conical surfaces, the apices of which meet at the two linear EPs $(\alpha,\beta,\delta)=(0,\pm\gamma,0)$. At large $\alpha$, the three–cusped shape remains but is deformed guaranteeing that the two cone-like surfaces originating at $\beta=\pm 1$ do not overlap; see Fig. 3.
Figure 3: Cross section of EP surface for fixed nonlinearity $\alpha$; axis scaled with scale factor $s=2(\sqrt{1+\alpha^{2}/\gamma^{2}}-1)$. Here, $\beta^{\prime}=\beta-\gamma$. Small $\alpha\ll\gamma$ (vicinity of linear EP, Eq. 5): deltoid or three-cusped hypocycloid broecker-lander.1975 ; poston-stewart.1978 ; saunders.80 of the elliptic umbilic surface with exact scaling $s=\alpha^{2}/\gamma^{2}$. Large $\alpha$ (farther away from linear EP, Eq. 4): exact scaling is lost, but three–cusped shape remains. The half-infinite line where PT symmetry is unbroken is indicated by the blue line.
In our study, the elements of the eigenvectors of the non-Hermitian eigenvalue problem $[M(\textbf{x})-\lambda]\textbf{x}=0$ are the state variables. Our strategy is then as follows: (more details are given in the SM): The eigenvalue equation constitutes two equations for the unkowns that we call state variables (and which determine the eigenvector elements; see below). We write those equations as gradients of an appropriately chosen Lyapunov potential, i.e. in the form ${\rm grad}\{V_{L}(\textbf{x})\}=0$. The condition for an EP (coalescing eigenvectors) is then the same as ${\rm det}\{h\}=0$. Finally, we perform a Taylor expansion of $V_{L}$ around the EP, allowing us to identify the Lyapunov potential with one of the elementary catastrophes, listed for example on p. 137 of saunders.80 and p. 11 of gilmore.81 , and reproduced in the SM. This then fixes the topological and geometric structure of the EP neighborhood. Below, we illustrate our approach for $2\times 2$ matrices with a second–order linear EP. In that case, we find that linear EPs are the organizing point of an elliptic umbilic singularity.
We write the nonlinear eigenvalue problem (now using the symbol $H$ as commonly used in the context of non-Hermitian Hamiltonians) as
$H\begin{bmatrix}x\\ y\end{bmatrix}=\begin{bmatrix}a&\beta\\ \beta&-a\end{bmatrix}\begin{bmatrix}x\\ y\end{bmatrix}=E\begin{bmatrix}x\\ y\end{bmatrix}\,,$ (1)
where $a=\delta+i\gamma+\alpha\left(|x|^{2}-|y|^{2}\right)$, $\alpha,\beta,\gamma,\delta\in\mathbb{R}$ are parameters, and all variables and parameters are assumed scaled and unitless. This parametric form is chosen for ease of analysis; the restrictions can be relaxed and the method is still applicable (see SM for details). In the linear limit, the two states are separated by $2\delta$ in frequency, and one state gains and the other dissipates at equal rates $|\gamma|$. The eigenvector, normalized to $1$, is written as $[x\,\,y]^{T}=(1/\textit{A})[1\,\,\tilde{w}]^{T}$, $\tilde{w}=we^{i\theta}$, $\textit{A}=\sqrt{1+w^{2}}$. Eliminating the eigenvalue $E$ from Eq. (1) reduces the matrix equation to a nonlinear equation in the complex eigenvector element $\tilde{w}$, the real and imaginary parts of which are
$2\alpha\left[\frac{1-w^{2}}{1+w^{2}}\right]+2\delta+\beta\left[w-\frac{1}{w}\right]\cos\theta$ $=0\,,$ (2)
$2\gamma+\beta\left[w+\frac{1}{w}\right]\sin\theta$ $=0\,.$ (3)
Formulating the eigenvalue problem this way enables the map to the function singularity problem to be constructed. In this context, we call $(w,\theta)$ the state variables and $(\alpha,\beta,\gamma,\delta)$ the control parameters. The first step in the map is to construct a Lyapunov function $V_{L}(w,\theta,\alpha,\beta,\gamma,\delta)$ such that its critical point equations in state variable space, $\partial V_{L}/\partial w=0$, $\partial V_{L}/\partial\theta=0$, are the same as the eigenvector equations (2) and (3). It can be directly verified that the Lyapunov function
$V_{L}(w,\theta,\alpha,\beta,\gamma,\delta)$ $=$ $-2\alpha\ln\left[\frac{g}{2}\right]+2\delta\ln w$ (4)
$+\beta g\cos\theta-2\gamma\left[\theta-\frac{3\pi}{2}\right]\,,$
with $g=w+w^{-1}$, satisfies this requirement.
To see the connection with catastrophe theory, we need to evaluate $V_{L}$ in the vicinity of the EPs. Since $V_{L}$ only has terms that are linear in the control parameters, we can reduce the dimensions of the parameter space by rescaling by one of the parameters, which we choose to be $\gamma$. The parameter space for the EP analysis, $(\alpha/\gamma,\beta/\gamma,\delta/\gamma)$, is thus three-dimensional. In the linear limit ($\alpha/\gamma=0$), the matrix $H$ has two separate EPs (bifurcation points) in the $(\beta/\gamma,\delta/\gamma)$ plane at $(\beta/\gamma,\delta/\gamma)=(\pm 1,0)$. The coalescing eigenvector element (degenerate critical point) is $(w,\theta)=(1,3\pi/2)$ for $\beta/\gamma=1$ and $(w,\theta)=(1,\pi/2)$ for $\beta/\gamma=-1$. Since the critical point behavior around the two linear EPs is related by reflection symmetry about the plane $\beta=0$, it suffices to analyze one, $\beta/\gamma=1$, of the two. We choose, for convenience, the $(w,\theta)$-independent terms in Eq. (4) for $V_{L}$ to shift $V_{L}$ to zero at the $\beta/\gamma=1$ linear EP. We firstly consider the properties of the nonlinear EP set in the immediate neighborhood of this linear EP. We write $r=w-1$, $\phi=\theta-3\pi/2$, $\beta^{\prime}/\gamma=\beta/\gamma-1$ and denote by $V_{EU}(r,\phi,\alpha/\gamma,\beta^{\prime}/\gamma,\delta/\gamma)$ the lowest-order terms in an expansion in $(r,\phi,\alpha/\gamma,\beta^{\prime}/\gamma,\delta/\gamma)$ of $V_{L}$. Then, we find
$V_{EU}(r,\phi,\alpha/\gamma,\beta^{\prime}/\gamma,\delta/\gamma)$ (5)
[…]
The form of Eq. (5) matches that of a universal unfolding of the $D_{-4}$ singularity, also called elliptic umbilic singularity, around the linear EP. Its degenerate critical points and bifurcations, which are well understood, are obtained as simultaneous solutions to three equations: $\partial V_{EU}/\partial r=0$, $\partial V_{EU}/\partial\theta=0$, and $\det\{h\}=0$, where $h$ is the Hessian matrix of $V_{EU}$. The singularity set, plotted in Fig. 2, consists of a pair of two conical surfaces with deltoid cross sections, the apices of which meet at the linear EPs at $(\alpha,\beta^{\prime},\delta)=(0,0,0)$ and $(0,-2\gamma,0)$. There are four and two critical points in state variable space at each control parameter point inside and outside the cones, respectively; also see Fig. 3 below. The ribs of each cone are cusp points where three critical points inside the cone meet and emerge as one critical point outside the cone. Each point on the surface between two ribs is a fold point where two critical points inside the cone meet and annihilate each other, thus vanishing on the outside. This arrangement of the fold and cusp singularities is characteristic of the elliptic umbilic catastrophe ($D_{-4})$. We call the meeting point of the cones the ‘organizing point’ of the catastrophe. We show in Fig. 3 that the cone-like topological structure is not limited to the immediate neighborhood of the linear EP, but remains valid even at large nonlinearities (large $\alpha$).
Regarding PT symmetry hassan-etal.2015 ; miri-alu.2016 ; ge-elganainy.2016 ; kominis-etal.2017 ; teimourpour-etal.2017 ; xia-etal.2021 ; lee-etal.2024 ; klauck-etal.2025 ; bender-hook.2024 , we note that the Hamiltonian in Eq. (1) is PT-symmetric, i.e. $[H,\mathrm{PT}]=0$, with parity represented by the $\sigma_{1}$ Pauli matrix, if $\delta=0$. In that case, PT symmetry is unbroken (i.e. the eigenenergies are real and the eigenvectors are common to $H$ and PT) if $|\beta|>\sqrt{\alpha^{2}+\gamma^{2}}$. This region is indicated as thick blue line in Fig. 3.
The power of using catastrophe theory in the context of EPs can be seen from the following considerations. First, the information on the geometric structure of the nonlinear EPs provides immediate knowledge of the number of eigenvalues as a function of the parameters used. The potential $V_{EU}$ has four critical points inside the cone and two outside the cone; exemplary plots of $V_{EU}$ are shown in the SM. This implies that we have four eigenvalues inside the cone and two outside, and the number changes at the folds and cusps, which represent the location of the EPs. This is illustrated in Fig. 4, which shows the eigenvalues of $H$ along the straight lines in ($\delta,\beta^{\prime})$-parameter space as indicated in each row of subplots. At the apices of the cones, Fig. 2, the four critical points coalesce, showing that, depending on how the limit is performed, the linear EP can be viewed as the limiting case of a 4th-order nonlinear EP. A stability analysis of the solutions is planned as future work.
Figure 4: Complex eigenvalues (real and imaginary parts) of $H$ in Eq (1) at fixed, nonzero $\alpha/\gamma=1$ across lines in the $(\beta^{\prime},\delta)$ plane shown as insets. Regardless of the lines’ directions, we have two and four eigenvalues outside and inside the cone, respectively, in agreement with the number of values of critical points of the Lyapunov potential (see SM for examples of the potential landscape). At the smooth parts of the deltoid boundary (fold lines), two states are created/annihilated, and two states are unaffected by the boundary (a,b,e,f). Crossing the cusp, (c,d), three states coalesce at a 3rd-order nonlinear EP, denoted as EP3, and one state is unaffected.
The second aspect highlighting the power of catastrophe theory lies in the fact that the linear matrix fixes already the elliptic umbilic catastrophe, and a broad class of matrices extending our case given in Eq. 1 will have the same singularity structure. For example, we show in the SM that any diagonal linear or nonlinear perturbation that depends only on $w$ through a function $F(w)$ with $F(w)=0$ and $dF(w)/dw\neq 0$ at the linear EP is in this category. It also includes the case of higher-order nonlinearities and additional linear parameters on the diagonal matrix elements. Furthermore, relaxing the assumption of a symmetric matrix and/or adding linear parameters to the off-diagonal matrix elements is allowed. Higher-power (smooth) nonlinearities $F(w)$ may limit the validity of the exact elliptic umbilic structure to a finite region around the linear EP. For the class of $2\times 2$ matrices delineated in this paper, this local validity is always maintained.
For $N\times N$ matrices with $N>2$, the dimension of the state variable space increases with $N$, the set of linear EPs is expected to have more structure than that of the $N=2$ case. More complicated, higher-order singularities, e.g. parabolic umbilics or singularities outside of the elementary catastrophe set, may be involved, most of which have higher codimensions (minimum dimension of the parameter space to capture the full structure of the singularity set) than the elliptic umbilic. But even when a relatively complete analysis comparable to the $2\times 2$ case cannot be done, using the tools of catastrophe theory and bifurcation theory may still provide fruitful insight into the topological/geometric structure of the singularity set.
In summary, we have given a topological and geometric interpretation of EPs in linear and nonlinear physical systems, in the sense that they correspond to degenerate critical points of a (Lyapunov) potential function. Therefore, the powerful theorems of catastrophe theory can be used to identify universal features such as the shapes of the set of EPs (singularity set) as a function of the physical parameters. From this (topological) perspective, a second order linear EP can be viewed as a fourth-order nonlinear EP in an extended parameter space. Future extensions of this work may include higher-dimensional linear matrices (see, e.g., Fig. 2 in binder-kwong.2021 ) and the case where the nonlinearity couples states beyond the $2\times 2$ subset, essentially also increasing the dimension of the Hamiltonian. Besides rigorously paving the way for future design of nonlinear non-Hermitian systems with novel functionalities, the universal topology and singularity structure reported in the present study also opens up new possibilities for the targeted design of nonlinear EP sensors.
Acknowledgements.
The authors gratefully acknowledge financial support for the Arizona group from the US National Science Foundation (NSF) under Grant No. DMR-1839570, and, for the Paderborn groups, by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) through Grant No. 467358803 and the transregional collaborative research center TRR 142 (Projects A04 and C10, Grant No. 231447078).